The temperature of the Earth's crust increases by
Answer #1

rate of heat transfred, dQ/dt = k*A*dT/t

= 0.8*(4*pi*Re^2)*3.33/100

= 0.8*(4*pi*(6.38*10^6)^2)*3.33/100

= 1.3626*10^13 J/s

heat transported out in one day, Q = (dQ/dt)*dt

= 1.3626*10^13*24*60*60

= 1.18*10^18 J <<<<<<<<<-------------------Answer